OpenStudy (anonymous):

Simplify: (3a + 5b)^2

OpenStudy (anonymous):

9a=25b

OpenStudy (anonymous):

9a+25b*

OpenStudy (anonymous):

\[(a+b)^2= a^2+2ab+b^2

OpenStudy (anonymous):

\[(a+b)^2= a^2+2ab+b^2\]